angelcook2007:

The function f(x) = −x2 + 28x − 192 models the hourly profit, in dollars, a shop makes for selling sodas, where x is the number of sodas sold and f(x) is the amount of profit. Part A: Determine the vertex. What does this calculation mean in the context of the problem? (5 points) Part B: Determine the x-intercepts. What do these values mean in the context of the problem? (5 points)

xhikari:

this could help you graph the equation.. https://www.desmos.com/calculator

angelcook2007:

Ok, so I'm not good at math

angelcook2007:

so how would I graph it

xhikari:

i think the 2 after -x is supposed to be like lower than the number right? i don't think it lets u type that on here

angelcook2007:

the 2 after the x is supposed to be an exponent

xhikari:

ohhh

xhikari:

oh okay lemme graph that

angelcook2007:

\[f(x) = −x^2 + 28x − 192\]

xhikari:

okay so heres the graph so you know what I'm talking about https://www.desmos.com/calculator/4sg7vl6plb

xhikari:

do you know what a vertex is?

angelcook2007:

it's like the highest point in the graph isn't it?

xhikari:

well it can be the highest or the lowest depending on if its a minimum or maximum

xhikari:

in this case it is the highest which means its a maximum

angelcook2007:

Okay

xhikari:

so the vertex would be (14,4)...

angelcook2007:

where do you get (14, 4) from?

xhikari:

one sec ill show you

angelcook2007:

because on the graph it's showing this

angelcook2007:

1 attachment
xhikari:

1 attachment
xhikari:

the highest point is the vertex

xhikari:

if it opened upwards then the lowest point would be the vertex

angelcook2007:

I was looking at the wrong spot my bad

xhikari:

its okay dw

carmelle:

Would you mind if I explained how to get the answer to part (sorry to disturb, I just wanna help) @xhikari

angelcook2007:

I was looking where the arch went past the X axis on both sides

carmelle:

part A

carmelle:

*

xhikari:

@carmelle wrote:
Would you mind if I explained how to get the answer to part (sorry to disturb, I just wanna help) @xhikari
sure ig

xhikari:

when it goes past the x axis those are the 'zeros' aka x intercepts which is part b

angelcook2007:

Oh okay, so theres 2 x intercepts ?

xhikari:

yes

angelcook2007:

ima go get somethin to drink rq if you don't mind @_@

xhikari:

yea np

angelcook2007:

it's like 80 degrees outside and it's almost 7 in the afternoon @_@

angelcook2007:

90 degrees sorry

angelcook2007:

anyways so, theres 2 x intercepts per graph, but what happens to 192 in the equation?

xhikari:

i dont think we need that bc its just asking for the vertex, x-intercepts, and what those mean.

angelcook2007:

Oh okay, so 192 doesn't count towards Part A nor B?

xhikari:

i dont think so. so we know that the vertex is (14,4) and the x-intercepts are 12 and 16

carmelle:

im gonna use a new method to solve this problem instead, are u familiar with completing the square?

angelcook2007:

@carmelle wrote:
im gonna use a new method to solve this problem instead, are u familiar with completing the square?
No?

angelcook2007:

@xhikari wrote:
i dont think so. so we know that the vertex is (14,4) and the x-intercepts are 12 and 16
Okay

xhikari:

sorry my mum called me down

xhikari:

do you still need me orr??

angelcook2007:

yeah, and it's alr

xhikari:

okay so i told you what the vertex means (maximum number of hourly profit (someone correct if I'm wrong)) so you have part A and now for part b we need to determine what the x interceots represent

angelcook2007:

So part A is maximum number of hourly profit?

xhikari:

yes but also include the vertex itself which is (14,4)

carmelle:

ok, so lets use the formula \[x = -\frac{ b }{ 2a }\] In this case a = -1 b = 28 When we plug those numbers in we get...\[-\frac{ 28 }{ 2(-1) }\] What do you get when you solve this equation @angelcook2007

carmelle:

sorry I took so long :(

angelcook2007:

oh okay, so like "The vertex is (14, 4), and this shows the maximum number of hourly profit"?

angelcook2007:

@carmelle wrote:
sorry I took so long :(
it's alr mate

carmelle:

sooo... what do you get when you solve the equation?

angelcook2007:

sorry, my questioncove tab decided to die

carmelle:

@angelcook2007 wrote:
sorry, my questioncove tab decided to die
its alr

angelcook2007:

anyways

angelcook2007:

solve \[\frac{ -28 }{ 2 -1}\]

angelcook2007:

???

carmelle:

yea... basically this... try solving it :)

carmelle:

what did you get?

angelcook2007:

14?

carmelle:

That's correct, now you find your x-vaule for your vertex :D (14, ) Now to find the y-value, you just substitute the `x` in the problem with 14 then solve for `y` \[f(x) = −x^2 + 28x − 192\] \[y = −(14)^2 + 28(14) − 192\] Can you try `solving for y`?

angelcook2007:

2?

angelcook2007:

not 2

angelcook2007:

4?

angelcook2007:

idek where 2 came from lol

carmelle:

correct, the y-value would be 4. That means the vertex is `(14,4)`

carmelle:

does that make sense ?

angelcook2007:

me and @xhikari found the vertex

angelcook2007:

bro it's so hot outside lol

carmelle:

yea, but you have to explain how you got the answer, not just say that u went on desmos ;/

angelcook2007:

oh ok lol

carmelle:

And you have to use the equation to find the answer lol

angelcook2007:

i'm honestly tempted to go eat some ice ngl

angelcook2007:

how hot is it where you live lol

angelcook2007:

it's almost 90 at my house

carmelle:

very hot, but lets continue and not get off track lol, do you still need help with part b?

angelcook2007:

lol

angelcook2007:

The x intercepts are 12 and 16 but idk what they represent

carmelle:

Well, the x-intercepts basically means that the shop will make 0 dollars when they sell 12 and 16 sodas soooo sorry for taking so long, im studying for my finals tomorrow lol

carmelle:

Do you need help on finding the x-intercepts btw ?

angelcook2007:

it's alright

angelcook2007:

@carmelle wrote:
Do you need help on finding the x-intercepts btw ?
wym?

carmelle:

@angelcook2007 wrote:
@carmelle wrote:
Do you need help on finding the x-intercepts btw ?
wym?
nvm I think u got it :)

angelcook2007:

oh lol

angelcook2007:

I randomly got a medal btw, what does that mean?

carmelle:

@angelcook2007 wrote:
I randomly got a medal btw, what does that mean?
medals are given from people who think you did a good job answering a question

angelcook2007:

Oh

angelcook2007:

okay, thanks lol

angelcook2007:

so the x intercepts mean that when they sell 12 and 16 sodas, they make a $0 profit or somthing?

carmelle:

@angelcook2007 wrote:
so the x intercepts mean that when they sell 12 and 16 sodas, they make a $0 profit or somthing?
yup- that's correct

angelcook2007:

oh okay, thank you